i want the interpretation of t test and f test based on excel

2. You have been asked to develop a multiple regression model to predict the annual number of traffic deaths. The data file "traffic death rate" contains 10 years of annual data on the death rates per 100 mlion vehicle miles (y), total travel in billion vehicle miles (x1), and average speed in miles per hour of all vehicles (x2). Compute the multiple regression of y on xl and x2 and write a report discussing your findings
Y X1 X2 738.0 52.6 5.2 5.3 767.0 53.8 5.4 805.0 55.8 5.6 847.0 55.9 5.5 889.0 56.4 5.7 930.0 57.3 5.5 962.0 58.0 1016.0 59.0 5.4 5.2 1071.0 60.0 4.9 1121.0 59.2 LO LO LO
Answer #1

2.

SP 1y (X1-Mx1)2 (x2-M2)2 X1-Mx1 SP 1x2 X2-M2 SPx2y Y-My -4.2 -0.17 30.022 0.714 741.72 -176.6 31187.56 17.64 10.332 0.21 442.

Sum of X1 = 9146
Sum of X2 = 568
Sum of Y = 53.7
Mean X1 = 914.6
Mean X2 = 56.8
Mean Y = 5.37
Sum of squares (SSX1) = 150038.4
Sum of squares (SSX2) = 51.14
Sum of products (SPX1Y) = -91.12
Sum of products (SPX2Y) = -0.65
Sum of products (SPX1X2) = 2648.7

Regression Equation = ŷ = b1X1 + b2X2 + a

b1 = ((SPX1Y)*(SSX2)-(SPX1X2)*(SPX2Y)) / ((SSX1)*(SSX2)-(SPX1X2)*(SPX1X2)) = -2938.22/657352.09 = -0.00447

b2 = ((SPX2Y)*(SSX1)-(SPX1X2)*(SPX1Y)) / ((SSX1)*(SSX2)-(SPX1X2)*(SPX1X2)) = 143824.58/657352.09 = 0.21879

a = MY - b1MX1 - b2MX2 = 5.37 - (0*914.6) - (0.22*56.8) = -2.96943

ŷ = -0.00447X1 + 0.21879X2 - 2.96943